### v1

# UIShift: Enhancing VLM-based GUI Agents through Self-supervised Reinforcement Learning

Abstract
Training effective Vision Language Models (VLMs) for GUI agents typically relies on supervised fine-tuning (SFT) over large-scale annotated datasets, where the collection process is labor-intensive and error-prone. In this work, we propose a self-supervised inverse dynamics task to enable VLMs to learn from GUI transition pairs by inferring the action that caused that transition. This training task offers two advantages: (1) It enables VLMs to ignore variations unrelated to user actions (e.g., background refreshes, ads) and to focus on true affordances such as buttons and input fields within complex GUIs. (2) The training data can be easily obtained from existing GUI trajectories without requiring human annotation, and it can be easily scaled through automatic offline exploration. Using this training task, we propose UIShift, a framework for enhancing VLM-based GUI agents through self-supervised reinforcement learning (RL). With only 2K training samples sourced from existing datasets, two VLMs—Qwen2.5-VL-3B and Qwen2.5-VL-7B—trained with UIShift achieve competitive or superior performance on grounding tasks (ScreenSpot-series benchmarks) and GUI automation tasks (AndroidControl), compared to SFT baselines and GUI-specific models that explicitly elicit reasoning abilities during RL. Our findings suggest a potential direction for enhancing VLMs for GUI agents by leveraging more self-supervised training data in the future.
1 Introduction
Mobile GUI agents [9, 11, 20, 30, 35, 39] interpret natural language instructions and perform low-level actions (e.g., click, scroll) directly on smartphone screens. They can control diverse apps as a human would, improving accessibility for users who are visually impaired, elderly, or have their hands occupied. Breakthroughs of Vision Language Models (VLMs) [2, 4, 26, 36] have reshaped the design paradigm of mobile GUI agents, transitioning from handcrafted heuristics to learned, vision-grounded policies. However, when facing complex multi-step tasks, these models struggle to deliver satisfactory accuracy [6, 20, 21, 40]. A common approach to enhance these VLMs is to apply supervised fine-tuning (SFT) on annotated datasets, e.g., using pre-recorded GUI trajectories with human-annotated task instructions [14, 22]. Although effective, collecting task-annotated GUI trajectories remains labor-intensive and error-prone [7, 8, 14, 22]. For example, collecting the AndroidControl [14] dataset takes one year of paid annotation effort to produce 15,283 task demonstrations. The high cost of the data collection process makes this paradigm difficult to scale.
In this study, we aim to leverage the abundance of unlabeled GUI trajectories, which can be easily obtained and scaled out from existing datasets or through automated offline exploration, e.g., MobileViews [8] at a million scale. To unlock the potential of such data, we propose a self-supervised training task for GUI agents, named k-step UI Transition. The task extends inverse-dynamics modeling from robotics and biomechanics applications [3, 27, 38]—where a network predicts the control command that links two consecutive physical states—by treating screenshots as states and GUI actions as commands. Concretely, the training samples consist of pairs of screenshots and , where is obtained by executing actions starting from state . The VLM is asked to predict the initial action that transforms into . This training task removes the need for predefined instructions or labeled trajectories, where ground-truth actions internally exist in GUI transition sequences. Compared to single-screen tasks (e.g., predicting the action to be executed on a screen based on task instructions), UI Transition enhances visual perception via explicit state comparison, and enables the model to better align actions with visual differences.
To exploit UI Transition, we adopt Group Relative Policy Optimization (GRPO) [24] instead of SFT. Under SFT, a cross-entropy loss is applied to enforce a single ground-truth action per screenshot pair, which impairs its performance when downstream tasks take a screenshot plus an instruction. In contrast, GRPO optimizes over multiple candidates: for each pair, it samples a group of plausible actions, scores each with the predefined reward function, and ranks them using group-wise normalized advantages. This approach encourages exploration, maintains instruction-following capability, and enhances generalization to downstream tasks, such as GUI grounding and task automation.
We propose UIShift, a one-stage RL training framework that fine-tunes VLMs using GRPO on UI Transition data alone. We fine-tune Qwen2.5-VL-3B and 7B models on 1K/2K UI Transition pairs. To show that UI Transition is a more effective task under the same RL framework, we compare UIShift against two annotation-dependent baselines: Task Automation Low and Task Automation High, which take a single screenshot and a step-level or task-level instruction to predict the next action that either completes the step or advances the task. To isolate the impact of the GRPO itself, we include a SFT variant on the same UI Transition pairs, without any reasoning prompts. This allows us directly compare of GRPO versus cross-entropy training on identical data. To further explore internal factors in UI Transition, we vary two settings: (1) Reasoning prompts: three train-inference configurations (reasoning in both training and inference, reasoning only during training, no reasoning at either stage). (2) Transition distance : number of steps between the two UI states, with . All models are evaluated on five benchmarks: three for grounding (ScreenSpot [5], ScreenSpot-V2 [32], ScreenSpot-Pro [13]) and two for task automation (AndroidControl-Low/High [14]).
We summarize our key results as follows.
UIShift, trained solely on self-supervised UI Transition data, consistently outperforms annotation-dependent baselines across both grounding and task automation benchmarks. This suggests that UI Transition offers a more effective learning signal for developing GUI understanding and generalization capabilities in VLMs.
Reasoning is not necessary for GUI-related tasks. We observe that models trained and evaluated without reasoning prompts perform comparably to, or even better than, those trained with reasoning. This aligns with the results reported in Perception-R1 [37] and UI-R1 [19]. Notably, GRPO outperforms SFT under the same data constraints, highlighting its efficiency when training data is limited.
Larger improves long-horizon planning, but yields the most balanced performance. Although achieves the highest accuracy on AndroidControl-High, it degrades grounding performance. The best balance occurs at , so we train UIShift-7B-2K with . Without reasoning prompts during training and inference, UIShift-7B-2K achieves 87.81% on ScreenSpot. On the mobile subset, it reaches 96.7% (text) and 85.2% (icon) accuracy, outperforming all prior methods. On ScreenSpot-V2, it reaches 90.3% accuracy and surpasses UI-TARS-7B (89.5%) trained with 50B tokens, setting a new state-of-the-art. UIShift-7B-2K also delivers consistent gains on task automation benchmarks.
Our contributions are summarized as follows.
We propose UIShift, the first self-supervised reinforcement training framework that combines the -step UI Transition task with GRPO to fine-tune VLMs.
We explore the necessity of reasoning in static GUI-related tasks, finding that reasoning is not essential for either training or inference.
We train UIShift-7B-2K model using UIShift with reasoning-free configuration, and evaluates its performance across five GUI benchmarks. Results show that it achieves better performance than models trained on annotation-dependent tasks.
2 Related Work
2.1 Mobile GUI Agents
Recent progress in mobile GUI agents has been driven by VLMs trained via SFT on large-scale datasets. These models learn to map instructions to GUI actions using instruction-following tasks, making high-quality annotations essential. Despite the availability of diverse GUI datasets [7, 8, 14, 15, 18, 22, 28], the quantity of high-quality annotations remains insufficient for robust training and usually requires significant human effort to scale. Prior SFT pipelines often incorporate out-of-domain image-caption pairs [11, 29] and supplement training with web and desktop data to improve cross-platform generalization [5]. As a result, the overall scale of training data tends to be large: UI-TARS [20] trains on 50B tokens; Uground [9] uses 1.3M screenshots to train a visual grounding model; OS-Atlas [32] leverages 13M UI elements for grounding pretraining. MobileVLM [31] incorporates an action prediction task during pretraining; however, but limits it to one-step transitions and still relies on annotation-heavy fine-tuning for downstream generalization. In this study, we propose a self-supervised training paradigm that eliminates reliance on large-scale human-annotated data. By constructing a -step inverse dynamics task where VLMs infer the action leading from the current to a future UI state, our approach captures long-horizon GUI dynamics and enables scalable training on large, underutilized, and unlabeled GUI transition datasets.
2.2 Rule-based Reinforcement Learning
Rule-based RL has proven to be a promising alternative to SFT. GRPO [24] uses a reward model to score each response and computes group relative advantages instead of training a critic model, whose size is comparable to the policy model, thereby significantly reducing computational cost. Reinforcement Learning with Verifiable Rewards [12] further emphasize the use of verifiable answers to design reliable reward signals. DeepSeek-R1 [10] shows that simple format and accuracy rewards are sufficient to surpass the performance of instruction-tuned models. This paradigm is particularly well-suited to GUI agents, whose action space consists of discrete action types with structured parameters, making correctness easy to verify. Several recent efforts have applied GRPO to GUI tasks: UI-R1 [19] conducts one-stage RL on 136 step-instruction samples; GUI-R1 [33] also adopts a one-stage RL strategy on 3K task-instruction samples from five platforms; InfiGUI-R1 [17] employs a two stage SFT+RL pipeline and scales to 32K samples from both GUI and non-GUI domains. While these works demonstrate the effectiveness of rule-based RL for GUI agents, they still rely on annotated instructions and require reasoning during training and inference. Differing from existing approaches, UIShift adopts a self-supervised training paradigm. We directly fine-tune VLMs via one-stage RL using only 2K UI Transition samples, yet achieve competitive performance and strong generalization across platforms and tasks.
3 UIShift Framework
UIShift is a one-stage RL training framework designed to fine-tune GUI-specific VLMs using self-supervised UI Transition task. This section first presents GRPO and the reward design employed in UIShift, and then introduces the -step UI Transition task along with its underlying rationales.
3.1 Preliminary: GRPO
GRPO [24] is a resource-efficient alternative to Proximal Policy Optimization (PPO) [23], a popular actor-critic RL algorithm. Unlike PPO, which requires a separate critic model to estimate value functions, GRPO directly computes normalized group-wise advantages using rewards from a reward model, thereby eliminating the need for value estimation and reducing computational overhead. DeepSeek-R1 [10] further simplifies this setup by replacing the reward model with a rule-based function composed of format and accuracy criteria. Following this design, the GRPO objective optimized in our framework is:
Specifically, for each question in the training set, we sample a group of outputs from the old policy using high temperature decoding, and compute the group-wise relative advantage for each output. A clipped surrogate objective, along with a KL divergence regularizer toward the reference policy , is then used to update model parameters and ensure training stability.
3.2 Reward Design in UIShift
To assess the impact of reasoning in GUI-related tasks, we design two training configurations: reasoning-enabled and reasoning-free, depending on whether the model is required to perform chain-of-thought reasoning during training. In both settings, we apply a rule-based reward function composed of a format reward and an accuracy reward :
Format reward. We adopt the format rules from DeepSeek-R1 [10] to ensure that model outputs follow the expected structure. In the reasoning-enabled setting, the model is asked to place intermediate reasoning inside <think>...</think> and the final answer inside <answer>...</answer>. In the reasoning-free setting, only the <answer>...</answer> block is required. Outputs that conform the expected structure receive ; otherwise, .
Accuracy reward. We use a unified action space during training and inference, covering five action types: click, scroll, open_app, navigate_back, and input_text. Each action is represented as a structured JSON object containing an action_type and its associated parameters. Specifically, click uses a coordinate pair (x, y) indicating the center point of the target UI element, scroll requires a direction (up, down, right, or left), open_app and input_text take a string parameter for the app name and text to be entered, respectively, and navigate_back is parameter-free. Predicted actions are considered correct if the action type and all parameters match the ground truth. For non-click actions, if all parameters exactly match, the model receives ; otherwise, . For click, we adopt a spatially tolerant strategy: the prediction is considered correct if the coordinate (x, y) fall within the bounding box (bbox) of the target UI element. This reflects real-world GUI behavior, where clicking anywhere within a UI component yields the same effect. To implement this, we extract the bbox of the target element from the corresponding view hierarchy file based on the ground-truth click action in the first screenshot of each UI Transition pair. This bbox is then appended to the training data. During training, the model is prompted to output only the coordinate (x, y), which are then compared against the ground truth bbox. If the predicted coordinate falls within the bbox, the model receives ; otherwise, .
3.3 K-step UI Transition Task and Training
As shown in Fig. 1, each training sample in the -step UI Transition task comprises a pair of UI screenshots: the current UI state and a future state , obtained by performing actions to . The model receives these two screenshots without any extra instructions and is tasked with predicting the first action that transforms into . While VLMs can understand single GUI screens due to the inclusion of GUI data in pretraining, they cannot capture changes across states or plan action sequences. The -step UI Transition task addresses this by forcing the model to compare and , identify the semantic change, localize the the responsible region in , and then output a corresponding structured action. This mirrors inverse-dynamics modeling in robotics, where predicting the control signal that caused a state change leads to representations that capture both “what changed” and “how to change” in the environment. In the GUI context, when the model must focus on “what changed” across two UI states, it naturally learns to ignore irrelevant UI elements and attend to actionable regions. That should boost robustness when applying the fine-tuned model to zero-shot grounding (identify a widget based on a text query) or to task automation benchmarks where it must plan several actions in sequence.
We fine-tune VLMs on the -step UI Transition task using GRPO. For each input pair, the model samples candidate actions, each scored by a rule-based reward that combines format correctness and action accuracy . We compute group-wise normalized advantages and optimize using the GRPO objective (Eq. 1). GRPO fits the UI Transition task because it enables the model to first propose a variaty of plausible actions and then rank them, providing richer feedback than one-shot labels in SFT. This naturally encourages exploration, unlike standard supervised losses that require exact matching behavior. By combining UI Transition with GRPO, our method could easily scale without annotated data and avoids overfitting to a single best action.
4 Experiments
We evaluate two fundamental capabilities of VLMs in GUI agents—instruction mapping and GUI grounding—across five benchmarks. Our comparisons span three dimensions: (1) prior methods, (2) annotation-dependent tasks, and (3) SFT baselines. We further investigate whether chain-of-thought reasoning helps and explore how k-step targets in UI Transition influence performance.
4.1 Experimental Setup
Models and training configurations. Using the training pipeline proposed in §3 and the open-source RL framework, VLM-R1 [25], we apply full-parameter reinforcement fine-tuning to Qwen2.5-VL-3B/7B [2] on 8A800 GPUs for 8 epochs. Our best model, UIShift-7B-2K, completes training in 26 hours with 2K UI Transition samples. Additional hyperparameter details are listed in Appendix A.
Training tasks and data selection. In UIShift, we adopt the k-step UI Transition alone and explore four variants with . We include two annotation-dependent tasks for comparison: (1) Task Automation Low provides a screenshot and a step instruction (e.g., “Click on the share icon”), which specifies the exact UI action to take; (2) Task Automation High provides a screenshot and a task instruction (e.g., “Share the news article on Gmail”), which describes the final goal, requiring the model to infer the next action to advance this task. We construct all training tasks from AndroidControl [14], which provides UI trajectories with screenshots and instructions. To ensure GUI diversity, we select training samples from different episodes across all tasks.
Benchmarks and evaluation metrics. For GUI grounding, we evaluate models on three benchmarks: ScreenSpot [5], ScreenSpot-V2 [32], and ScreenSpot-Pro [13]. ScreenSpot and ScreenSpot-V2 provide evaluations across mobile, web, and desktop platforms, while ScreenSpot-Pro focuses on high-resolution desktop interfaces with fine-grained UI components. In all three benchmarks, the model receives a textual description and is required to localize a UI element by predicting its screen coordinate (x, y). A prediction is correct if the coordinate falls within the ground-truth bounding box. We report grounding accuracy across various UI types, platforms, and application categories.
For GUI task automation, we adopt AndroidControl-Low and AndroidControl-High [14], both of which provide a screenshot and an instruction, and require the model to predict the next action. Each predicted action consists of an action type and its associated parameters. While AndroidControl-Low includes an explicit step-level instruction as direct guidance, AndroidControl-High provides only a task-level instruction, requiring the model to infer the current progress within a multi-step task, thereby making the decision more challenging. We report three evaluation metrics: (i) action type accuracy (Type), the proportion of predicted action types that exactly match the ground truth; (ii) grounding accuracy (Grounding), the percentage of samples where the predicted coordinates fall within the ground-truth bounding box, computed over samples where both the ground-truth and predicted actions are click; (iii) success rate (SR), the proportion of predictions where both the action type and all associated parameters exactly match the ground truth.
Reasoning configurations and comparative setup. To study the effectiveness of chain-of-thought reasoning in VLMs, we experiment with three different configurations during training and inference: (1) both training and inference with reasoning prompts (fully reasoning-dependent); (2) training with reasoning but inference without it (inference reasoning-free); (3) training and inference both without reasoning (fully reasoning-free). To further assess the effectiveness of RL, we also train SFT baselines on the same three datasets—UI Transition, Task Automation Low, Task Automation High—under the fully reasoning-free setting.
4.2 Results on GUI Grounding and Task Automation
| AndroidControl-Low | AndroidControl-High | |||||||
| Model | Method | Data | Type | GR | SR | Type | GR | SR |
| Qwen2.5-VL-7B [2] | Zero Shot | - | 89.3 | 68.3 | 65.4 | 70.3 | 47.6 | 37.2 |
| UI Transition with SFT | ||||||||
| UIShift-7B-SFT () | SFT | 1K | 96.8 | 23.3 | 38.9 | 70.8 | 15.2 | 20.4 |
| UI Transition with GRPO (UIShift, self-supervised) for | ||||||||
| UIShift-7B-1K () | 1K | 90.9 | 88.4 | 79.8 | 81.2 | 65.0 | 54.2 | |
| UIShift-7B-2K () | 2K | 97.5 | 88.3 | 86.0 | 82.4 | 65.1 | 55.2 | |
| UIShift-7B-1K () | 1K | 96.9 | 88.4 | 85.7 | 80.6 | 67.4 | 55.7 | |
| UIShift-7B-1K () | 1K | 97.3 | 87.7 | 85.4 | 80.0 | 66.3 | 54.5 | |
| UIShift-7B-1K () | RL | 1K | 97.4 | 88.4 | 85.9 | 81.7 | 66.8 | 56.1 |
| Instruction-Dependent Training | ||||||||
| OS-Atlas-7B [32] | SFT | 2.3M | 93.6 | 88.0 | 85.2 | 85.2 | 78.5 | 71.2 |
| Aguvis-7B [34] | SFT | 1M | - | - | 80.5 | - | - | 61.5 |
| UI-TARS-7B [20] | SFT | 50B tokens | 98.0 | 89.3 | 90.8 | 83.7 | 80.5 | 72.5 |
| UI-R1-3B [19] | RL | 136 | 94.3 | 82.6 | - | - | - | - |
| GUI-R1-7B [33] | RL | 3K | - | - | - | 71.6 | 65.6 | 51.7 |
| InfiGUI-R1-3B [17] | SFT+RL | 12K + 20K | 96.0 | 93.2 | 92.1 | 82.7 | 74.4 | 71.1 |
GUI task automation benchmarks. As shown in Tab. 1, UIShift-7B consistently improves over Qwen2.5-VL-7B baselines on both AndroidControl-Low and High. Using only 1K samples, UIShift-7B-1K achieves 79.8% on AndroidControl-Low and 54.2% on High, improving by 14.4% and 17%, respectively. Scaling to 2K samples further improves results to 86.0% and 55.2%, indicating that even small-scale UI Transition training benefits downstream automation tasks.
| Mobile | Desktop | Web | |||||||
| Model | Method | Data | Text | Icon | Text | Icon | Text | Icon | Avg. |
| Qwen2.5-VL-7B [2] | Zero Shot | - | 97.1 | 81.2 | 86.6 | 70.0 | 87.4 | 78.6 | 84.9 |
| UI Transition with GRPO (UIShift, self-supervised) for | |||||||||
| UIShift-7B-1K () | 1K | 96.7 | 85.2 | 87.6 | 73.6 | 89.6 | 81.1 | 86.9 | |
| UIShift-7B-2K () | 2K | 96.7 | 85.2 | 91.2 | 75.7 | 89.6 | 82.0 | 87.8 | |
| UIShift-7B-1K () | 1K | 96.7 | 85.2 | 87.6 | 74.3 | 88.7 | 80.1 | 86.6 | |
| UIShift-7B-1K () | 1K | 96.7 | 83.0 | 89.2 | 73.6 | 89.6 | 80.6 | 86.6 | |
| UIShift-7B-1K () | RL | 1K | 96.7 | 85.2 | 86.6 | 75.0 | 88.7 | 80.1 | 86.6 |
| Instruction-Dependent Training | |||||||||
| CogAgent-18B [11] | SFT | - | 67.0 | 24.0 | 74.2 | 20.0 | 70.4 | 28.6 | 47.4 |
| SeeClick-9.6B [5] | SFT | 1M | 78.0 | 52.0 | 72.2 | 30.0 | 55.7 | 32.5 | 53.4 |
| UGround-7B [9] | SFT | 1.3M | 82.8 | 60.3 | 82.5 | 63.6 | 80.4 | 70.4 | 73.3 |
| OS-Atlas-7B [32] | SFT | 2.3M | 93.0 | 72.9 | 91.8 | 62.9 | 90.9 | 74.3 | 82.5 |
| ShowUI-2B [16] | SFT | 256K | 92.3 | 75.5 | 76.3 | 61.1 | 81.7 | 63.6 | 75.1 |
| Aguvis-7B [34] | SFT | 1M | 95.6 | 77.7 | 93.8 | 67.1 | 88.3 | 75.2 | 84.4 |
| UI-TARS-7B [20] | SFT | 50B tokens | 94.5 | 85.2 | 95.9 | 85.7 | 90.0 | 83.5 | 89.5 |
| UI-R1-3B [19] | RL | 136 | 95.6 | 84.7 | 90.2 | 59.3 | 85.2 | 73.3 | 83.3 |
| UI-R1-E-3B [19] | RL | 2K | 97.1 | 83.0 | 95.4 | 77.9 | 91.7 | 85.0 | 89.2 |
| GUI-R1-7B [33] | RL | 3K | - | - | 91.8 | 73.6 | 91.3 | 75.7 | - |
| InfiGUI-R1-3B [17] | SFT+RL | 12K + 20K | 97.1 | 81.2 | 94.3 | 77.1 | 91.7 | 77.6 | 87.5 |
| Mobile | Desktop | Web | |||||||
| Model | Method | Data | Text | Icon | Text | Icon | Text | Icon | Avg. |
| Qwen2.5-VL-7B [2] | Zero Shot | - | 98.6 | 87.2 | 93.8 | 88.7 | 90.2 | 81.3 | 87.9 |
| UI Transition with GRPO (UIShift, self-supervised) for | |||||||||
| UIShift-7B-1K () | 1K | 99.0 | 89.1 | 90.7 | 75.7 | 91.0 | 83.3 | 89.5 | |
| UIShift-7B-2K () | 2K | 99.3 | 89.6 | 92.3 | 77.9 | 92.3 | 82.3 | 90.3 | |
| UIShift-7B-1K () | 1K | 99.0 | 89.1 | 90.7 | 75.7 | 92.3 | 81.3 | 89.5 | |
| UIShift-7B-1K () | 1K | 98.6 | 87.7 | 91.2 | 75.7 | 92.3 | 81.8 | 89.3 | |
| UIShift-7B-1K () | RL | 1K | 98.6 | 90.5 | 90.2 | 76.4 | 91.5 | 83.3 | 89.8 |
| Instruction-Dependent Training | |||||||||
| SeeClick-9.6B [5] | SFT | 1M | 78.4 | 50.7 | 70.1 | 29.3 | 55.2 | 32.5 | 55.1 |
| OS-Atlas-7B [32] | SFT | 2.3M | 95.2 | 75.8 | 90.7 | 63.6 | 90.6 | 77.3 | 84.1 |
| UI-TARS-7B [20] | SFT | 50B tokens | 96.9 | 89.1 | 95.4 | 85.0 | 93.6 | 85.2 | 91.6 |
| UI-R1-3B [19] | RL | 136 | 96.2 | 84.3 | 92.3 | 63.6 | 89.2 | 75.4 | 85.4 |
| UI-R1-E-3B [19] | RL | 2K | 98.2 | 83.9 | 94.8 | 75.0 | 93.2 | 83.7 | 89.5 |
GUI grounding benchmarks. UIShift-7B-2K achieves strong grounding accuracy on ScreenSpot and ScreenSpot-V2, as shown in Tab. 2 and Tab. 3. On mobile, it achieves 96.7% for text and 85.2% for icon on ScreenSpot, and 99.3% for text and 89.6% for icon on ScreenSpot-V2. In terms of average accuracy across all three platforms, UIShift-7B-2K achieves 87.8% on ScreenSpot and 90.3% on ScreenSpot-V2, closely matching UI-TARS-7B (89.5% and 91.6%, respectively), which is trained on roughly 50B tokens. On the high-resolution ScreenSpot-Pro benchmark, UIShift demonstrates improved grounding accuracy compared to the Qwen2.5-VL-7B baseline. Detailed results of ScreenSpot-Pro are shown in Appendix. B.
4.3 Comparing Reasoning Configurations
| Model | Reasoning? | Mobile | Desktop | Web | Avg. | ||||
| Train | Infer | Text | Icon | Text | Icon | Text | Icon | ||
| UIShift-3B | Y | Y | 94.5 | 73.8 | 89.2 | 64.3 | 85.2 | 69.4 | 80.9 |
| Y | N | 94.1 | 74.2 | 84.5 | 61.4 | 82.2 | 66.0 | 78.8 | |
| N | N | 95.2 | 74.7 | 86.6 | 62.1 | 82.2 | 66.5 | 79.6 | |
| UIShift-7B | Y | Y | 97.1 | 82.1 | 89.7 | 72.1 | 91.3 | 79.1 | 86.6 |
| Y | N | 97.1 | 83.8 | 84.5 | 73.6 | 89.1 | 80.6 | 86.1 | |
| N | N | 96.7 | 85.2 | 87.6 | 73.6 | 89.6 | 81.1 | 86.9 | |
| Reasoning? | AndroidControl-Low | AndroidControl-High | |||||||
| Model | Method | Train | Infer | Type | GR | SR | Type | GR | SR |
| SFT | N | N | 95.6 | 9.4 | 28.0 | 67.0 | 7.5 | 11.3 | |
| Y | Y | 93.1 | 87.6 | 82.1 | 79.6 | 59.9 | 50.1 | ||
| Y | N | 91.2 | 88.7 | 80.7 | 78.2 | 60.9 | 49.7 | ||
| UIShift-3B | RL | N | N | 96.9 | 88.6 | 85.4 | 78.8 | 60.8 | 50.0 |
| SFT | N | N | 96.8 | 23.3 | 38.9 | 70.8 | 15.2 | 20.4 | |
| Y | Y | 94.0 | 88.6 | 83.1 | 75.7 | 71.0 | 54.7 | ||
| Y | N | 97.3 | 88.3 | 85.7 | 81.1 | 65.2 | 54.3 | ||
| N | N | 90.9 | 88.4 | 79.8 | 81.7 | 65.0 | 54.2 | ||
| UIShift-7B | RL | N | N | 97.4 | 88.4 | 85.9 | 81.7 | 66.8 | 56.1 |
We compare three reasoning configurations in GRPO-based training—fully reasoning-dependent, inference reasoning-free, and fully reasoning-free—across AndroidControl-Low, AndroidControl-High, and ScreenSpot. As shown in Tab. 4 and Tab. 5, for each model size, the three reasoning configurations achieve comparable performance on AndroidControl-High and ScreenSpot. On AndroidControl-Low, the fully reasoning-free configuration performs best: the 3B model reaches its peak at , while the 7B achieves the highest accuracy at . These results suggest that the presence of reasoning during inference has minimal impact. Additionally, as shown in Tab. 5, GRPO-trained models still outperform their SFT-trained counterparts under fully reasoning-free mode. Given the similar performance across reasoning configurations and the substantial training time difference (13 hours for reasoning-free vs. 19 hours for reasoning-dependent on the 7B model with 1K samples), we conclude that reasoning-free training offers an efficient and effective strategy without compromising generalization performance.
4.4 Comparing Training Tasks
| Reasoning? | Mobile | Desktop | Web | |||||||
| Size | Task | Train | Infer | Text | Icon | Text | Icon | Text | Icon | Avg. |
| Y | N | 94.5 | 73.8 | 84.5 | 61.4 | 80.4 | 62.1 | 77.8 | ||
| Task Auto. (Low) | Y | Y | 95.6 | 74.7 | 89.7 | 64.3 | 85.2 | 68.0 | 81.1 | |
| Y | N | 94.5 | 73.4 | 85.6 | 61.4 | 80.9 | 64.6 | 78.4 | ||
| Task Auto. (High) | Y | Y | 95.2 | 75.1 | 90.2 | 65.7 | 85.7 | 70.9 | 81.9 | |
| Y | Y | 94.5 | 73.8 | 85.1 | 89.2 | 85.2 | 69.4 | 80.9 | ||
| Y | N | 94.1 | 74.2 | 84.5 | 61.4 | 82.2 | 66.0 | 78.8 | ||
| 3B | UI Transition () | N | N | 95.2 | 74.7 | 86.6 | 62.1 | 82.2 | 66.5 | 79.6 |
| Y | N | 96.7 | 83.0 | 87.6 | 70.7 | 89.6 | 79.6 | 85.9 | ||
| Task Auto. (Low) | Y | Y | 97.1 | 80.8 | 90.7 | 71.4 | 90.9 | 77.7 | 86.1 | |
| Y | N | 96.7 | 85.2 | 86.1 | 73.6 | 88.3 | 80.6 | 86.3 | ||
| Task Auto. (High) | Y | Y | 96.3 | 82.1 | 90.2 | 75.0 | 90.4 | 77.2 | 86.3 | |
| Y | Y | 97.1 | 82.1 | 90.2 | 89.7 | 91.3 | 79.1 | 86.6 | ||
| Y | N | 97.1 | 83.8 | 84.5 | 73.6 | 89.1 | 80.6 | 86.1 | ||
| 7B | UI Transition () | N | N | 96.7 | 85.2 | 87.6 | 73.6 | 89.6 | 81.1 | 86.9 |
| AndroidControl-Low | AndroidControl-High | |||||
| Size | Training Task | Reasoning? (Train) | Infer: N | Infer: Y | Infer: N | Infer: Y |
| Task Auto. (Low) | Y | 85.2 | 85.4 | 34.6 | 35.5 | |
| Task Auto. (High) | Y | 84.4 | 84.8 | 49.3 | 49.4 | |
| UI Transition () | Y | 80.7 | 82.1 | 49.7 | 50.1 | |
| 3B | UI Transition () | N | 85.4 | - | 50.0 | - |
| Task Auto. (Low) | Y | 85.5 | 86.4 | 47.6 | 48.3 | |
| Task Auto. (High) | Y | 83.8 | 80.8 | 55.4 | 56.1 | |
| UI Transition () | Y | 85.7 | 83.1 | 54.3 | 54.7 | |
| UI Transition () | N | 79.8 | - | 54.2 | - | |
| 7B | UI Transition () | N | 85.9 | - | 56.1 | - |
We compare the performance of models trained on UI Transition with those trained on annotation-dependent tasks—Task Automation Low and Task Automation High—across AndroidControl and ScreenSpot benchmarks. As shown in Tab. 6, all three training tasks achieve comparable grounding performance on ScreenSpot for both 3B and 7B models. Notably, the 1-step UI Transition variant under the fully reasoning-free configuration slightly outperforms the instruction-based baselines. As shown in Tab. 7, annotation-dependent tasks perform well on their respective in-domain benchmarks: Task Automation Low on AndroidControl-Low and Task Automation High on AndroidControl-High. However, we observe limited generalization when models trained on the simpler Task Automation Low are evaluated on the more challenging AndroidControl-High, which requires broader task-level understanding. In contrast, UI Transition exhibits strong and balanced performance across two task automation tasks, matching or even surpassing the annotation-dependent baselines. We attribute this to the task formulation: the model receives a current state and a future state as a purely visual target. This allows the model to focus on global screen-level comparison and perception. Compared to Task Automation Low and High, where the model must align the current UI screen with a textual instruction representing either a 1-step or k-step goal, UI Transition provides a more direct visual comparison between UI states and thus eliminates the need for screen-text alignment. These results highlight that UI Transition delivers strong performance without extra annotation, making it an effective and scalable alternative for training GUI-specific models.
4.5 Effectiveness of K in K-step UI Transition
The k-step UI Transition takes the current UI state and a future state , and trains the model to predict the initial action in the path from to , i.e., the action leading to . By adjusting k, this formulation accommodates both step-level and task-level prediction. We evaluate how different values of k affect model performance across five benchmarks using 1K training samples, and additionally assess the effect of increasing the data size to 2K for . As shown in Tab. 1, with 1K training samples, achieves the highest accuracy on AndroidControl-Low and AndroidControl-High, but shows a slight performance drop on ScreenSpot-Pro (Appendix. B). Meanwhile, leads to more stable performance across all grounding benchmarks. When the data size increases to 2K with , performance improves consistently across all benchmarks. Notably, accuracy on AndroidControl-Low increases by 6.2%. We attribute this improvement to the alignment between the formulation of the 1-step UI Transition and the nature of AndroidControl-Low, as each step instruction corresponds to a one-step textual target. These results indicate that the choice of is a critical factor and should be aligned with the granularity and structure of the target task.
5 Conclusion
Collecting task-annotated GUI trajectories is costly and error-prone, which constrains the scalability of supervised fine-tuning on instructional data. In this work, we propose UIShift, a self-supervised, one-stage RL framework that fine-tunes GUI-specific VLMs on the -step UI Transition task using GRPO. By predicting the first action that moves UI state to given the pair (, ), the model learns to map screen-level visual differences to low-level actions. This objective lets it generalizes to GUI grounding and task automation without extra labels, eliminates annotation costs, and unlocks large-scale, unlabeled GUI transition datasets. Without reasoning prompts at either training or inference time, UIShift-7B-2K outperforms annotation-dependent baselines on five static benchmarks, showing that UI Transition provides a strong learning signal for GUI task generalization.
This study has two main limitations. First, we train on 1K/2K samples due to compute constraints; accessing how performance scales with larger datasets is the next step. Second, our evaluation is confined to static benchmarks; future work will extend to dynamic ones such as AndroidWorld [21].
Appendix A Training Hyperparameters
We use a fixed set of hyperparameters across all tasks to ensure a fair comparison. Details are provided in Table 8.
| Hyperparameter | Value |
| learning_rate | from 1e-6 to 0 |
| temperature | 0.9 |
| num_generations | 8 |
| num_train_epochs | 8 |
| max_prompt_length | 1024 |
| max_completion_length | 256 |
| per_device_train_batch_size | 2 |
| gradient_accumulation_steps | 8 |
| (clipping parameter) | 0.2 |
| (KL coefficient) | 0.04 |
Appendix B Evaluation Results
| CAD | Development | Creative | Scientific | Office | OS | ||||||||
| Model | Text | Icon | Text | Icon | Text | Icon | Text | Icon | Text | Icon | Text | Icon | Avg. |
| Zero Shot | |||||||||||||
| Qwen2.5-VL-7B [2] | 18.8 | 1.6 | 48.1 | 6.9 | 26.7 | 6.3 | 52.1 | 5.5 | 48.6 | 17 | 33.6 | 13.5 | 25.8 |
| UI Transition with GRPO (UIShift, self-supervised) for | |||||||||||||
| UIShift-7B-1K () | 21.8 | 1.6 | 50.7 | 9.0 | 31.3 | 7.0 | 51.4 | 5.5 | 48.0 | 18.9 | 41.1 | 14.6 | 27.8 |
| UIShift-7B-2K () | 21.8 | 3.1 | 49.4 | 6.9 | 35.9 | 6.3 | 52.1 | 6.4 | 52.5 | 20.8 | 37.4 | 10.1 | 28.2 |
| UIShift-7B-1K () | 21.8 | 1.6 | 50.7 | 6.9 | 30.3 | 7.0 | 50.7 | 4.6 | 49.7 | 20.8 | 37.4 | 15.7 | 27.4 |
| UIShift-7B-1K () | 19.8 | 0.0 | 48.7 | 7.6 | 29.8 | 7.7 | 52.8 | 4.6 | 52.0 | 22.6 | 38.3 | 10.1 | 27.2 |
| UIShift-7B-1K () | 20.3 | 3.1 | 50.0 | 6.9 | 27.3 | 5.6 | 49.3 | 6.4 | 44.1 | 15.1 | 37.4 | 12.4 | 25.7 |
| Instruction-Dependent Training | |||||||||||||
| CogAgent-18B [11] | 7.1 | 3.1 | 14.9 | 0.7 | 9.6 | 0.0 | 22.2 | 1.8 | 13.0 | 0.0 | 5.6 | 0.0 | 7.7 |
| SeeClick-9.6B [5] | 2.5 | 0.0 | 0.6 | 0.0 | 1.0 | 0.0 | 3.5 | 0.0 | 1.0 | 0.0 | 2.8 | 0.0 | 1.1 |
| OS-Atlas-7B [32] | 12.2 | 4.7 | 33.1 | 1.4 | 28.8 | 2.8 | 37.5 | 7.3 | 33.9 | 5.7 | 27.1 | 4.5 | 18.9 |
| ShowUI-2B [16] | 2.5 | 0.0 | 16.9 | 1.4 | 9.1 | 0.0 | 13.2 | 7.3 | 15.3 | 7.5 | 10.3 | 2.2 | 7.7 |
| UGround-7B [9] | 14.2 | 1.6 | 26.6 | 2.1 | 27.3 | 2.8 | 31.9 | 2.7 | 31.6 | 11.3 | 17.8 | 0.0 | 16.5 |
| UI-TARS-7B [20] | 20.8 | 9.4 | 58.4 | 12.4 | 50.0 | 9.1 | 63.9 | 31.8 | 63.3 | 20.8 | 30.8 | 16.9 | 35.7 |
| UI-R1-3B [19] | 11.2 | 6.3 | 22.7 | 4.1 | 27.3 | 3.5 | 42.4 | 11.8 | 32.2 | 11.3 | 13.1 | 4.5 | 17.8 |
| UI-R1-E-3B [19] | 37.1 | 12.5 | 46.1 | 6.9 | 41.9 | 4.2 | 56.9 | 21.8 | 65.0 | 26.4 | 32.7 | 10.1 | 33.5 |
| GUI-R1-7B [33] | 23.9 | 6.3 | 49.4 | 4.8 | 38.9 | 8.4 | 55.6 | 11.8 | 58.7 | 26.4 | 42.1 | 16.9 | - |
| InfiGUI-R1-3B [17] | 33.0 | 14.1 | 51.3 | 12.4 | 44.9 | 7.0 | 58.3 | 20.0 | 65.5 | 28.3 | 43.9 | 12.4 | 35.7 |
Both training and evaluation of UIShift models are conducted without reasoning prompts. On ScreenSpot-Pro, UIShift-7B-2K achieves an average grounding accuracy of 28.2%, improving 2.4% over the untuned Qwen2.5-VL-7B baseline (25.8%).